# Cosmological constraints from the EFT power spectrum and tree-level bispectrum of 21cm intensity maps

### v1

@@ 4 Results
Figure 3: Constraints on $\Lambda$CDM cosmological parameters from the combination of the EFT power spectrum and tree-level bispectrum of a HIRAX-like survey for $k_{\max,0}=0.25h/$Mpc in Equation 2.24. All three cases presented here consider Planck priors (see section 3) and utilise the signal from the whole redshift range.
@@ 4.1 $\Lambda$CDM
We also find that the 1-loop power spectrum exhibits different correlations between parameters than the bispectrum. This indicates the possibility to break degeneracies and tightly constrain these parameters. The significant advantages in combining the two correlators is not evident here, due to the relatively small values considered for $k_{\max}^{B}(z)$ in the tree-level bispectrum analysis, compared to those used for the power spectrum EFT.
Figure 4: As in Figure 3, for the $\gamma{\rm CDM}\,$model.
@@ 4.5 Distance and growth rate errors
The constraints are derived after marginalising over the nuisance parameters for both correlators, while the remaining model parameters are considered free (Equation 3.6). Figure 6 (upper panels) presents the results for the marginalized relative errors as a function of redshift.
@@ 4.6 Bias parameters
Figure 6: Upper panels: Forecasts of marginalized relative errors on $D_{A}$, $H$, and $f$ from the power spectrum (red), bispectrum (blue), and their combination (orange), in the pessimistic case (Equation 2.25). Lower panels: Similarly, for the bias parameters in Equation 2.3.
@@ 5 Conclusions
In order to remain within the domain of validity of the tree-level bispectrum and 1-loop EFT power spectrum, we have imposed appropriate scale cuts, as illustrated in Equation 2.22Equation 2.25. For the EFT power spectrum, our baseline is a ‘pessimistic’ $k^{P}_{\rm max}$ cut, with errors also computed for a more optimistic choice.
@@
Compared to the tree-level power spectrum from our previous work [20], we find that EFT model improves precision on $\Omega_{\rm b}$, $\Omega_{\rm c}$, and $h$, due to the additional information from small scales, as probed by the 1-loop power spectrum. However, this is not the case for $A_{s}$ and $n_{s}$, which are less sensitive to this regime. The comparison with [20] also demonstrates the potential of the EFT power spectrum in breaking parameter degeneracies. This is clearly seen in the contour plots of Figure 3Figure 5.

### v2

@@ 4 Results
Figure 4: As in Figure 3, for the $\gamma{\rm CDM}\,$model.
@@ 4.1 $\Lambda$CDM
The 1-loop power spectrum and the tree-level bispectrum have different dependencies on the bias and cosmological parameters. This indicates the possibility to break degeneracies and improve constraints, which we intend to explore in follow-up work. The significant advantages in combining the two correlators is not evident here, due to the relatively small values considered for $k_{\max}^{B}(z)$ in the tree-level bispectrum analysis, compared to those used for the power spectrum EFT.
@@ 4.4 Summary of results on $\Lambda$CDM, $\gamma$CDM and $w_{0}w_{a}$
Figure 6: Upper panels: Forecasts of marginalized relative errors on $D_{A}$, $H$, and $f$ from the power spectrum (red), bispectrum (blue), and their combination (orange), in the pessimistic case (Equation 25). Lower panels: Similarly, for the bias parameters in Equation 3.
@@ 4.5 Distance and growth rate errors
The constraints are derived after marginalising over the nuisance parameters for both correlators, while the remaining model parameters are considered free (Equation 31). Figure 6 (upper panels) presents the results for the marginalized relative errors as a function of redshift.
@@ 5 Conclusions
In order to remain within the domain of validity of the tree-level bispectrum and 1-loop EFT power spectrum, we have imposed appropriate scale cuts, as illustrated in Equation 22Equation 25. For the EFT power spectrum, our baseline is a ‘pessimistic’ $k^{P}_{\rm max}$ cut, with errors also computed for a more optimistic choice.
@@
Previous work that uses the combination of the power spectrum and bispectrum in cosmological constraints is mainly focused on galaxy surveys (see e.g. [61] and references cited therein). In the case of 21cm intensity mapping, there are cosmological forecasts from the 1-loop power spectrum [15, 16], but we are not aware of other work that combines the 1-loop power spectrum with the bispectrum. For the 21cm tree-level power spectrum + bispectrum, the papers [18, 19] investigate future constraints on primordial non-Guassianity. Compared to the tree-level power spectrum + bispectrum forecasts from our previous work [20], we find that the EFT model improves precision on $\Omega_{\rm b}$, $\Omega_{\rm c}$, and $h$, due to the additional information from small scales, as probed by the 1-loop power spectrum. However, this is not the case for $A_{s}$ and $n_{s}$, which are less sensitive to this regime. The comparison with [20] also demonstrates the potential of the EFT power spectrum in breaking parameter degeneracies. This is clearly seen in the contour plots of Figure 3Figure 5.